OpenStudy (anonymous):

ryan puts $500 into a bank account. the bank pays 5% compound interest per year. how much is the interest after one year?

Parth (parthkohli):

Continuous compounding. \( \color{Black}{\Rightarrow 500(1 + 0.05) \large^n }\)

Parth (parthkohli):

\(n\textbf{ is the number of years.}\)

OpenStudy (anonymous):

ok

Parth (parthkohli):

\( \color{Black}{\Rightarrow 500(1.05)^1 }\) \( \color{Black}{\Rightarrow 500 \times 1.05 }\)

OpenStudy (anonymous):

$525?

OpenStudy (anonymous):

is it?

Parth (parthkohli):

This is compounded interest.

OpenStudy (anonymous):

im confused :s

Parth (parthkohli):

And yes it is $525 :)

OpenStudy (anonymous):

oh ok there's part 2

OpenStudy (anonymous):

lol sorry parth read mistake ;)

Parth (parthkohli):

Interest = Amount - Principal Interest = 525 - 500 Interest = 25

OpenStudy (anonymous):

work out the total amount he has in his bank account after 2 yrs

Parth (parthkohli):

@ZhangYan if it's one year then both simple and compound may apply :) so your answer is correct but not the method :S

OpenStudy (anonymous):

work out the total amount he has in his bank account after 2 yrs